Can you please help me to Solve this equation containing a rational exponent on the variable.
4x^3/4-31=77

Answer by mananth(16949)

4x^3/4 -31=77
4x^3/4=77+31
4x^3/4 =108
take the 4th power of the equation
(4x^(3/4))^4= 108)^4
256x^3=108^4
x^3=108^4/256
x^3=531441
x= cuberoot(531441)
x=81
m.ananth@hotmail.ca